# Twenty years of Nešetřil’s classification programme of Ramsey classes

### v3

@@ 7.2.4 Tree of types
An enumerated structure is simply a structure $\mathbf{A}$ with underlying set $A=|A|=\{0,1,\ldots|A|-1\}$. Fix a countably infinite enumerated structure $\mathbf{A}$. Given vertices $u,v$ and an integer $n$ satisfying $\min(u,v)\geq n\geq 0$, we write $u\sim^{\mathbf{A}}_{n}v$ and say that $u$ and $v$ have the same (quantifier-free) type over $\{0,1,\ldots,n-1\}$, if the structures induced by $\mathbf{A}$ on $\{0,1,\ldots,n-1,u\}$ and $\{0,1,\ldots,n-1,v\}$ are isomorphic via the map which is the identity on $\{0,...,n-1\}$ and sends $u$ to $v$. We write $u]^{\mathbf{A}}_{n}$ for the $\sim^{\mathbf{A}}_{n}$-equivalence class of vertex $u$.
@@
Let $\mathbf{A}$ be a countably infinite (relational) enumerated structure. Given $n<\omega$, write $\mathbb{T}_{\mathbf{A}}(n)=\omega/\!\sim^{\mathbf{A}}_{n}$. A (quantifier-free) 1-type is any member of the disjoint union $\mathbb{T}_{\mathbf{A}}:=\bigsqcup_{n<\omega}\mathbb{T}_{\mathbf{A}}(n)$. We turn $\mathbb{T}_{\mathbf{A}}$ into a tree as follows. Given $x\in\mathbb{T}_{\mathbf{A}}(m)$ and $y\in\mathbb{T}_{\mathbf{A}}(n)$, we declare that $x\leq^{\mathbb{T}}_{\mathbf{A}}y$ if and only if $m\leq n$ and $x\supseteq y$.
The tree of 1-types of a given structure is useful for giving upper bounds on big Ramsey degrees, since all the existing proofs are in fact structured as Ramsey-type results about colouring subtrees of the tree of 1-types (or a related object). In the other direction, they are also useful for constructing unavoidable colourings and for characterising big Ramsey degrees.
It is interesting to notice that while seeking a Ramsey expansion of a given class is usually a process that needs a good intuition or a good guess, determining characterisations of big Ramsey degrees can be done more systematically by analysing subtrees of the tree of types and comparing them with the upper bound theorems. This method was implicit in the early characterisations of big Ramsey degrees [188, 57, 16] and is explicitly applied in [14]. In fact, one can also use it for discovering canonical Ramsey expansions of amalgamation classes. As a proof of concept, Hubička used this method to rediscover that the canonical Ramsey expansion of triangle-free graphs are free orderings, while the canonical Ramsey expansion of posets are linear extensions [148].
@@ 7.2.5 Big Ramsey structures
    The Rado graph, and more generally, unconstrained structures in binary languages (this is an easy consequence of the characterisation of big Ramsey degrees by Laflamme, Sauer, and Vuksanovic [188]).
@@ 7.3 Generalizations of the KPT-correspondence and Ramsey classes
A categorical approach to Fraïssé limits was developed by Kubiś [175], leading to a notion of weak Ramsey property and an extension of KPT [12]. See also related results by Mašulović [195].
@@ 8.1.1 $H_{4}$-free 3-hypertournaments
One can then ask which subsets of $\{C_{4},O_{4},H_{4}\}$ one can forbid in order to obtain a Fraïssé class. A simple Ramsey argument shows that every sufficiently large 3-hypertournament contains a copy of $C_{4}$; and all the four remaining subsets of $\{C_{4},O_{4},H_{4}\}$ give rise to a Fraïssé class. Particularly interesting are two of them: The $O_{4}$-free (or even) 3-hypertournaments are, when ordered, first-order interdefinable with ordered two-graphs (see Example 5). More precisely, restricting the 1-to-1 correspondence between linearly ordered 3-hypertournaments and linearly ordered 3-uniform hypergraphs to the $O_{4}$-free case, one gets the class of linearly ordered two-graphs. Consequently, their canonical Ramsey expansion also adds a linear order and a graph from the switching class.
@@
The class of all $H_{4}$-free 3-hypertournaments is, at this point, at the very top of the list of these challenging examples. As opposed to other, longer-standing open concrete classes, this one is relatively recent (and thus has seen much fewer person-hours spent trying to solve it), it is a strong amalgamation class in a finite relational language and has an extremely simple description. See [166, Chapter 2] for a discussion of possible ways how things might resolve based on the answer to Question 8.2.
@@ 8.1.4 Finite measure algebras and equipartitions
Giordano and Pestov [110] proved that the automorphism group of every standard probability space is extremely amenable. In [174], Kechris, Sokić, and Todorčević identified a structural Ramsey statement which would imply this result: A finite dyadic measure algebra is a finite structure $\mathbf{A}=(A,\wedge,\vee,0,1,\mu)$, where $(A,\wedge,\vee,0,1)$ is a finite Boolean algebra and $\mu\colon A\to[0,1]$ is a positive measure with values in the dyadic rationals. The structural Ramsey statement identified in [174] says that the class of all naturally ordered finite dyadic measure algebras is Ramsey (see [174] for a definition of a natural ordering). This is, in turn, equivalent to the Ramsey property for finite Boolean algebras with antilexicographic orderings (see Example 6) where one only considers maps where every atom is sent to the join of the same number of atoms, which can also be seen as a version of the dual Ramsey theorem for equipartitions (see [174]; they call it the homogeneous dual Ramsey theorem). To our best knowledge, all these Ramsey-theoretic statements are still open.
@@ 8.1.5 Finite groups
Note that while there are some partition theorems for finite abelian groups [295], they do not seem to be structural, and thus not relevant for this question. We weakly conjecture that the answer is negative as we expect Hall’s universal locally finite group to have a complex structure of imaginaries. But even if the answer happens to positive, it feels even harder to prove it using Theorem 6.7 than, for example, Boolean algebras (see Example 6).
@@ 8.2 EPPA
Note that the $\{H_{4},O_{4}\}$-free 3-hypertournaments do not have EPPA as these structures become just linear orders after naming a least vertex. We expect that the $O_{4}$-free 3-hypertournaments will behave as a combination of two-graphs (see [87] and Example 5) and 3-hypertournaments with no restrictions.
@@
Two-graphs (see Example 5) turned out to be a very interesting example for EPPA [87]. They are one of the reducts (in the sense of Section 6.4.1) of the random graph, and one can ask about EPPA of the other reducts:
@@
There are five reducts. The random graph itself and the set with no structure have been solved. For two-graphs, coherence is open. The last two reducts are the complementing graph (adding an isomorphism between the random graph and its complement) and the “complementing two-graph” (the join of the automorphism group of the generic two-graph and the complementing graph) and for them both EPPA and coherent EPPA are open. In his Bachelor thesis, Beliayeu [21] proved EPPA for the class of all finite graphs with loops with complementing embeddings (that is, every vertex may, but does not have to, have a loop, and loops also get complemented). This created a particularly interesting situation: It is open whether graphs with complementing embeddings have EPPA, but there are two possible expansions of graphs with complementing embeddings which have EPPA: Graphs with normal embeddings, and graphs with loops with complementing embeddings. Moreover, both of them can be seen as reducts of the countable random graphs with a generic unary relation, and they are incomparable. It would be interesting to see if the meet of these two expansions are graphs with complementing embeddings, and whether the meet has EPPA.
@@
Does the class of all finite partial Steiner triple systems have EPPA, where one only wants to extend partial automorphism between closed substructures? (A sub-hypergraph $H$ of a Steiner triple system $S$ is closed if whenever $\{x,y,z\}$ is a triple of $S$ and $x,y\in H$, then $z\in H$.)
@@
In [87] it is proved that this class has EPPA. In fact, the automorphism group of the generic two-graph does have a dense locally-finite subgroup (a dynamical consequence of coherent EPPA, see Section 5 of [87]), but the EPPA-witnesses from [87] are not coherent. Two-graphs are thus a concrete candidate for an answer to a question of Siniora whether there is a class of finite structures with EPPA but not coherent EPPA [265, Question 3].
@@ 8.3 Big Ramsey structures
Therefore, before one tries to identify meaningful conjectures, more concrete basic examples should be looked at and solved. In fact, the existence of a big Ramsey structure (and the exact characterisation of big Ramsey degrees) has not been proved for any structure in a language with a relation of arity greater than 2. In particular, we have the following problem:
@@
Given a relational language $L$ with finitely many relational symbols of each arity, characterize the big Ramsey degrees for the class $\mathcal{K}$ of all finite $L$-structures where tuples in each relations are injective. Does the Fraïssé limit of $\mathcal{K}$ admit a big Ramsey structure?
@@
When free amalgamation classes are understood, one can start aiming for a big Ramsey analogue of Theorems 6.6 or 6.7 (or some other variant which can handle similar structures). In the Ramsey classes world, various variants of metric spaces turned out to be useful examples to look at first. Theorem 7.17 implies that homogeneous metric spaces with finitely many distances have finite big Ramsey degrees; the bounds on them produced by Theorem 7.17 are very far from optimal though. One should attempt to characterize the exact big Ramsey degrees for these metric spaces, and if possible, find the corresponding big Ramsey structures, in the hope of uncovering the core of the combinatorial nature of these structures.
@@
Let $S\subseteq\mathbb{R}^{{>}0}$ be finite and let $\mathbb{U}_{S}$ be a countable metric spaces with distances from $S$ universal for all countable metric spaces with distances from $S$. Characterise the big Ramsey degrees of $\mathbb{U}_{S}$. Does it admit a big Ramsey structure?
@@
We have seen that the class of finite Boolean algebras with antilexicographic orderings is Ramsey and that it is an easy consequence of the dual Ramsey theorem (see Example 6). In the infinite setting, these two statements diverge. While the Carlson–Simpson theorem [70] is a natural infinitary extension of the dual Ramsey theorem, one may also study the big Ramsey degrees of the countable atomless Boolean algebra (i. e., the Fraïssé limit of the class of all finite Boolean algebras). Shortly before finishing this survey it was shown that the big Ramsey degrees are in fact infinite for the subalgebra with 3 atoms [13], and it turns out that this is related to continuous colourings of 3-elements subsets of $\mathbb{Q}$ and topological copies [288, Theorem 6.33]. However we can still ask:
@@ Appendix A Partite construction
From the induced partite construction, two other constructions can be derived. In some cases, performing the induced partite construction once does not solve the problem, but after performing it multiple times, magic happens and problems disappear. Intuitively, the $\langle\mathbf{A},\mathbf{B},\mathbf{C}\rangle$-hypergraph (the hypergraph of copies in the Ramsey witness, see Section 6.2) becomes more and more sparse with each iteration of the partite construction. This iterated form is discussed in Section A.3 where we prove Theorem 6.6 which then implies, for example, that ordered metric spaces are Ramsey (Theorem 6.5). The last discussed extension is done by replacing one of the two key steps of the partite construction by another partite construction. We call this the recursive form and discuss one application of it in Section A.4 to prove an extension of the unrestricted Nešetřil–Rödl theorem for structures in languages with functions. Combining this with the iterated partite construction then yields the main results of [141] (Theorems 6.4 and 6.7) in their full strength.
@@ A.1 Non-induced partite construction
$L_{P}$-partite systems are a generalization of multipartite graphs. Note that the second condition allows a single vertex to appear multiple times in $\vec{x}$. Given an $L_{P}$-partite system $\mathbf{A}$ and a vertex $v\in A$ we denote by $\pi_{\mathbf{A}}(v)$ the unique predicate $p$ containing $v$. This defines a function $\pi_{\mathbf{A}}\colon A\to L_{P}\setminus P$ which we call projection. We denote by $A_{p}$ the set of all vertices of $\mathbf{A}$ with projection $p$ and call it the partition $p$.
@@ A.2.1 Examples of applications
Fix $L$ and $\mathcal{F}$ as in the statement. Let $\mathbf{A}$ and $\mathbf{B}$ be finite ordered $\mathcal{F}$-free structures. Use Theorem 2.5 to obtain an ordered $L$-structure $\mathbf{D}$ satisfying $\mathbf{D}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$, and then apply Theorem A.6 to get a $\mathbf{D}$-partite system $\mathbf{P}$ such that $\mathbf{P}|_{L}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$. Point (3) of Theorem A.6 implies that every irreducible structure of $\mathbf{P}|_{L}$ embeds to $\mathbf{B}$, and consequently $\mathbf{P}|_{L}$ is $\mathcal{F}$-free, as every $\mathbf{F}\in\mathcal{F}$ is irreducible, and hence if there was an embedding $\mathbf{F}\to\mathbf{\mathbf{missing}}P|_{L}$ then there would also be an embedding $\mathbf{F}\to\mathbf{B}$, a contradiction.
@@
Let $L$ contain two binary relations << and $\prec$, and let $\mathbf{A}$ and $\mathbf{B}$ be finite $L$-structures such that $(A,\prec_{\mathbf{A}})$ and $(B,\prec_{\mathbf{B}})$ are partial orders and $(A,<_{\mathbf{A}})$ and $(B,<_{\mathbf{B}})$ their respective linear extensions. Use Theorem 2.5 to obtain an ordered $L$-structure $\mathbf{D}$ satisfying $\mathbf{D}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$. Notice that while we know that $(\mathbf{D},<_{\mathbf{D}})$ is a linear order, we do not know much about $(\mathbf{D},\prec_{\mathbf{D}})$.
@@
Let $\mathbf{P}_{i}$ be a picture and denote by $\prec^{\prime}_{\mathbf{P}_{i}}$ the transitive closure of $\prec_{\mathbf{P}_{i}}$. Then
@@
    If $(u,v)\in\prec_{\mathbf{P}_{i}}$ then $(u,v)\in<_{\mathbf{P}_{i}}$. Formally $\prec_{\mathbf{P}_{i}}\subseteq<_{\mathbf{P}_{i}}$. Notice that neither $\prec_{\mathbf{P}_{i}}$ nor $_{\mathbf{P}_{i}}$ are orders.
@@
    If $(u,v)\in\prec^{\prime}_{\mathbf{P}_{i}}\setminus\prec_{\mathbf{P}_{i}}$ then $(u,v)\notin<_{\mathbf{P}_{i}}$. (Or in other words, $\prec^{\prime}_{\mathbf{P}_{i}}\cap<_{\mathbf{P}_{i}}\subseteq\prec_{\mathbf{P}_{i}}$.)
Invariant 1 is easy to check for $\mathbf{P}_{0}$, since $\prec_{\mathbf{P}_{0}}$ is transitive and consequently $\prec^{\prime}_{\mathbf{P}_{0}}=\prec_{\mathbf{P}_{0}}$. Assume that $\mathbf{P}_{i}$, for $i<m$ satisfies Invariant 1. Clearly, $\mathbf{P}_{i+1}|_{\alpha_{i}(\mathbf{A})}$ also satisfies Invariant 1 as it has homomorphism-embedding to $\mathbf{A}$ where $\prec_{\mathbf{A}}$ is a partial order and $_{\mathbf{A}}$ its linear extension. It remains to notice that Invariant 1 is preserved by free amalgamations.
Note that $_{\mathbf{P}}$ has a homomorphism-embedding to $_{\mathbf{D}}$ which is a linear order. Consequently, $_{\mathbf{P}}$ is acyclic. The first point of Invariant 1 implies that $\prec_{\mathbf{P}}$ is also acyclic. It remains to construct an $L$-structure $\mathbf{C}$ by putting $C=P$, letting $_{\mathbf{C}}$ extend $_{\mathbf{P}}$ to a linear order, and putting $\prec_{\mathbf{C}}$ to be the transitive closure of $\prec_{\mathbf{P}}$. By the second point of Invariant 1 we know that every embedding $f\colon\mathbf{B}\to\mathbf{P}|_{L}$ is also an embedding $f\colon\mathbf{B}\to\mathbf{C}$. ∎

### v5

@@ 7.2.4 Tree of types
An enumerated structure is simply a structure $\mathbf{A}$ with underlying set $A=|A|=\{0,1,\ldots|A|-1\}$. Fix a countably infinite enumerated structure $\mathbf{A}$. Given vertices $u,v$ and an integer $n$ satisfying $\min(u,v)\geq n\geq 0$, we write $u\sim^{\mathbf{A}}_{n}v$ and say that $u$ and $v$ have the same (quantifier-free) type over $\{0,1,\ldots,n-1\}$, if the structures induced by $\mathbf{A}$ on $\{0,1,\ldots,n-1,u\}$ and $\{0,1,\ldots,n-1,v\}$ are isomorphic via the map which is the identity on $\{0,\ldots,n-1\}$ and sends $u$ to $v$. We write $u]^{\mathbf{A}}_{n}$ for the $\sim^{\mathbf{A}}_{n}$-equivalence class of vertex $u$.
@@
Let $\mathbf{A}$ be a countably infinite (relational) enumerated structure. Given $n<\omega$, write $\mathbb{T}_{\mathbf{A}}(n)=\omega/\sim^{\mathbf{A}}_{n}$. A (quantifier-free) 1-type is any member of the disjoint union $\mathbb{T}_{\mathbf{A}}\coloneq\bigsqcup_{n<\omega}\mathbb{T}_{\mathbf{A}}(n)$. We turn $\mathbb{T}_{\mathbf{A}}$ into a tree as follows. Given $x\in\mathbb{T}_{\mathbf{A}}(m)$ and $y\in\mathbb{T}_{\mathbf{A}}(n)$, we declare that $x\leq^{\mathbb{T}}_{\mathbf{A}}y$ if and only if $m\leq n$ and $x\supseteq y$.
The tree of 1-types of a given structure is useful for giving upper bounds on big Ramsey degrees, since all the existing proofs are in fact structured as Ramsey-type results about colouring subtrees of the tree of 1-types (or a related object). In the other direction, they are also useful for constructing unavoidable colourings and for characterizing big Ramsey degrees.
It is interesting to notice that while seeking a Ramsey expansion of a given class is usually a process that needs a good intuition or a good guess, determining characterizations of big Ramsey degrees can be done more systematically by analysing subtrees of the tree of types and comparing them with the upper bound theorems. This method was implicit in the early characterizations of big Ramsey degrees [193, 57, 15] and is explicitly applied in [16]. In fact, one can also use it for discovering canonical Ramsey expansions of amalgamation classes. As a proof of concept, Hubička used this method to rediscover that the canonical Ramsey expansion of triangle-free graphs are free orderings, while the canonical Ramsey expansion of posets are linear extensions [153].
@@ 7.2.5 Big Ramsey structures
    The Rado graph, and more generally, unconstrained structures in binary languages (this is an easy consequence of the characterization of big Ramsey degrees by Laflamme, Sauer, and Vuksanovic [193]).
@@ 7.3 Generalizations of the KPT-correspondence and Ramsey classes
A categorical approach to Fraïssé limits was developed by Droste and Gübel [75], and later further extended by Kubiś [180], leading to a notion of weak Ramsey property and an extension of KPT [12]. See also related results by Mašulović [199].
@@ 8.1.1 $H_{4}$-free 3-hypertournaments
One can then ask which subsets of $\{C_{4},O_{4},H_{4}\}$ one can forbid in order to obtain a Fraïssé class. A simple Ramsey argument shows that every sufficiently large 3-hypertournament contains a copy of $C_{4}$; and all the four remaining subsets of $\{C_{4},O_{4},H_{4}\}$ give rise to a Fraïssé class. Particularly interesting are two of them: The $O_{4}$-free (or even) 3-hypertournaments are, when ordered, first-order interdefinable with ordered two-graphs (see Example 6.28). More precisely, restricting the 1-to-1 correspondence between linearly ordered 3-hypertournaments and linearly ordered 3-uniform hypergraphs to the $O_{4}$-free case, one gets the class of linearly ordered two-graphs. Consequently, their canonical Ramsey expansion also adds a linear order and a graph from the switching class.
@@
The class of all $H_{4}$-free 3-hypertournaments is, at this point, at the very top of the list of these challenging examples. As opposed to other, longer-standing open concrete classes, this one is relatively recent (and thus has seen much fewer person-hours spent trying to solve it), it is a strong amalgamation class in a finite relational language and has an extremely simple description. See [171, Chapter 2] for a discussion of possible ways how things might resolve based on the answer to Question 8.2.
@@ 8.1.4 Finite measure algebras and equipartitions
Giordano and Pestov [114] proved that the automorphism group of every standard probability space is extremely amenable. In [179], Kechris, Sokić, and Todorčević identified a structural Ramsey statement which would imply this result: A finite dyadic measure algebra is a finite structure $\mathbf{A}=(A,\wedge,\vee,0,1,\mu)$, where $(A,\wedge,\vee,0,1)$ is a finite Boolean algebra and $\mu\colon A\to[0,1]$ is a positive measure with values in the dyadic rationals. The structural Ramsey statement identified in [179] says that the class of all naturally ordered finite dyadic measure algebras is Ramsey (see [179] for a definition of a natural ordering). This is, in turn, equivalent to the Ramsey property for finite Boolean algebras with antilexicographic orderings (see Example 6.29) where one only considers maps where every atom is sent to the join of the same number of atoms, which can also be seen as a version of the dual Ramsey theorem for equipartitions (see [179]; they call it the homogeneous dual Ramsey theorem). To our best knowledge, all these Ramsey-theoretic statements are still open.
@@ 8.1.5 Finite groups
Note that while there are some partition theorems for finite abelian groups [299], they do not seem to be structural, and thus not relevant for this question. We weakly conjecture that the answer is negative as we expect Hall’s universal locally finite group to have a complex structure of imaginaries. But even if the answer happens to be positive, it feels even harder to prove it using Theorem 6.15 than, for example, Boolean algebras (see Example 6.29).
@@ 8.2 EPPA
Note that the $\{H_{4},O_{4}\}$-free 3-hypertournaments do not have EPPA as these structures become just linear orders after naming a least vertex. We expect that the $O_{4}$-free 3-hypertournaments will behave as a combination of two-graphs (see [91] and Example 6.28) and 3-hypertournaments with no restrictions.
@@
Two-graphs (see Example 6.28) turned out to be a very interesting example for EPPA [91]. They are one of the reducts (in the sense of Section 6.4.1) of the random graph, and one can ask about EPPA of the other reducts:
@@
There are five reducts. The random graph itself and the set with no structure have been solved. For two-graphs, coherence is open. The last two reducts are the complementing graph (adding an isomorphism between the random graph and its complement) and the “complementing two-graph” (the join of the automorphism group of the generic two-graph and the complementing graph) and for them both EPPA and coherent EPPA are open. In his Bachelor thesis, Beliayeu [21] proved EPPA for the class of all finite graphs with loops with complementing embeddings (that is, every vertex may, but does not have to, have a loop, and loops also get complemented). This created a particularly interesting situation: It is open whether graphs with complementing embeddings have EPPA, but there are two possible expansions of graphs with complementing embeddings which have EPPA: Graphs with normal embeddings, and graphs with loops with complementing embeddings. Moreover, both of them can be seen as reducts of the countable random graph with a generic unary relation, and they are incomparable. It would be interesting to see if the meet of these two expansions are graphs with complementing embeddings, and whether the meet has EPPA.
@@
Does the class of all finite partial Steiner triple systems have EPPA, where one only wants to extend partial automorphism between closed substructures? (A sub-hypergraph $H$ of a partial Steiner triple system $S$ is closed if whenever $\{x,y,z\}$ is a triple of $S$ and $x,y\in H$, then $z\in H$.)
@@
In [91] it is proved that this class has EPPA. In fact, the automorphism group of the generic two-graph does have a dense locally-finite subgroup (a dynamical consequence of coherent EPPA, see Section 5 of [91]), but the EPPA-witnesses from [91] are not coherent. Two-graphs are thus a concrete candidate for an answer to a question of Siniora whether there is a class of finite structures with EPPA but not coherent EPPA [269, Question 3].
@@ 8.3 Big Ramsey structures
Therefore, before one tries to identify meaningful conjectures, more concrete basic examples should be looked at and solved. In fact, the existence of a big Ramsey structure (and the exact characterization of big Ramsey degrees) has not been proved for any structure in a language with a relation of arity greater than 2. In particular, we have the following problem:
@@
Given a relational language $L$ with finitely many relational symbols of each arity, characterize the big Ramsey degrees for the class $\mathcal{K}$ of all finite $L$-structures where tuples in each relation are injective. Does the Fraïssé limit of $\mathcal{K}$ admit a big Ramsey structure?
@@
When free amalgamation classes are understood, one can start aiming for a big Ramsey analogue of Theorems 6.11 or 6.15 (or some other variant which can handle similar structures). In the Ramsey classes world, various variants of metric spaces turned out to be useful examples to look at first. Theorem 7.24 implies that homogeneous metric spaces with finitely many distances have finite big Ramsey degrees; the bounds on them produced by Theorem 7.24 are very far from optimal though. One should attempt to characterize the exact big Ramsey degrees for these metric spaces, and if possible, find the corresponding big Ramsey structures, in the hope of uncovering the core of the combinatorial nature of these structures.
@@
Let $S\subseteq\mathbb{R}^{{>}0}$ be finite and let $\mathbb{U}_{S}$ be a countable metric spaces with distances from $S$ universal for all countable metric spaces with distances from $S$. Characterize the big Ramsey degrees of $\mathbb{U}_{S}$. Does it admit a big Ramsey structure?
@@
We have seen that the class of finite Boolean algebras with antilexicographic orderings is Ramsey and that it is an easy consequence of the dual Ramsey theorem (see Example 6.29). In the infinite setting, these two statements diverge. While the Carlson–Simpson theorem [70] is a natural infinitary extension of the dual Ramsey theorem, one may also study the big Ramsey degrees of the countable atomless Boolean algebra (i. e., the Fraïssé limit of the class of all finite Boolean algebras). Shortly before finishing this survey it was shown that the big Ramsey degrees are in fact infinite for the subalgebra with 3 atoms [13], and it turns out that this is related to continuous colourings of 3-elements subsets of $\mathbb{Q}$ and topological copies [292, Theorem 6.33]. However we can still ask:
@@ Appendix A Partite construction
From the induced partite construction, two other constructions can be derived. In some cases, performing the induced partite construction once does not solve the problem, but after performing it multiple times, magic happens and problems disappear. Intuitively, the $\langle\mathbf{A},\mathbf{B},\mathbf{C}\rangle$-hypergraph (the hypergraph of copies in the Ramsey witness, see Section 6.2) becomes more and more sparse with each iteration of the partite construction. This iterated form is discussed in Section A.3 where we prove Theorem 6.11 which then implies, for example, that ordered metric spaces are Ramsey (Theorem 6.7). The last discussed extension is done by replacing one of the two key steps of the partite construction by another partite construction. We call this the recursive form and discuss one application of it in Section A.4 to prove an extension of the unrestricted Nešetřil–Rödl theorem for structures in languages with functions. Combining this with the iterated partite construction then yields the main results of [146] (Theorems 6.5 and 6.15) in their full strength.
@@ A.1 Non-induced partite construction
$L_{P}$-partite systems are a generalization of multipartite graphs. Note that the second condition allows a single vertex to appear multiple times in $\vec{x}$. Given an $L_{P}$-partite system $\mathbf{A}$ and a vertex $v\in A$ we denote by $\pi_{\mathbf{A}}(v)$ the unique predicate $p$ containing $v$. This defines a function $\pi_{\mathbf{A}}\colon A\to L_{P}\setminus L$ which we call projection. We denote by $A_{p}$ the set of all vertices of $\mathbf{A}$ with projection $p$ and call it the partition $p$.
@@ A.2.1 Examples of applications
Fix $L$ and $\mathcal{F}$ as in the statement. Let $\mathbf{A}$ and $\mathbf{B}$ be finite ordered $\mathcal{F}$-free structures. Use Theorem 2.7 to obtain an ordered $L$-structure $\mathbf{D}$ satisfying $\mathbf{D}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$, and then apply Theorem A.10 to get a $\mathbf{D}$-partite system $\mathbf{P}$ such that $\mathbf{P}|_{L}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$. Point (3) of Theorem A.10 implies that every irreducible structure of $\mathbf{P}|_{L}$ embeds to $\mathbf{B}$, and consequently $\mathbf{P}|_{L}$ is $\mathcal{F}$-free, as every $\mathbf{F}\in\mathcal{F}$ is irreducible, and hence if there was an embedding $\mathbf{F}\to\mathbf{\mathbf{}}P|_{L}$ then there would also be an embedding $\mathbf{F}\to\mathbf{B}$, a contradiction.
@@
Let $L$ contain two binary relations << and $\ll$, and let $\mathbf{A}$ and $\mathbf{B}$ be finite $L$-structures such that $(A,\ll_{\mathbf{A}})$ and $(B,\ll_{\mathbf{B}})$ are partial orders and $(A,<_{\mathbf{A}})$ and $(B,<_{\mathbf{B}})$ their respective linear extensions. Use Theorem 2.7 to obtain an ordered $L$-structure $\mathbf{D}$ satisfying $\mathbf{D}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$. Notice that while we know that $(\mathbf{D},<_{\mathbf{D}})$ is a linear order, we do not know much about $(\mathbf{D},\ll_{\mathbf{D}})$.
@@
Let $\mathbf{P}_{i}$ be a picture and denote by $\ll^{\prime}_{\mathbf{P}_{i}}$ the transitive closure of $\ll_{\mathbf{P}_{i}}$. Then
@@
    If $(u,v)\in{\ll_{\mathbf{P}_{i}}}$ then $(u,v)\in{<_{\mathbf{P}_{i}}}$. Formally ${\ll_{\mathbf{P}_{i}}}\subseteq{<_{\mathbf{P}_{i}}}$. Notice that neither $\ll_{\mathbf{P}_{i}}$ nor $_{\mathbf{P}_{i}}$ are orders.
@@
    If $(u,v)\in{\ll^{\prime}_{\mathbf{P}_{i}}}\setminus{\ll_{\mathbf{P}_{i}}}$ then $(u,v)\notin{<_{\mathbf{P}_{i}}}$. (Or in other words, ${\ll^{\prime}_{\mathbf{P}_{i}}}\cap{<_{\mathbf{P}_{i}}}\subseteq{\ll_{\mathbf{P}_{i}}}$.)
Invariant 1 is easy to check for $\mathbf{P}_{0}$, since $\ll_{\mathbf{P}_{0}}$ is transitive and consequently $\ll^{\prime}_{\mathbf{P}_{0}}=\ll_{\mathbf{P}_{0}}$. Assume that $\mathbf{P}_{i}$, for $i<m$ satisfies Invariant 1. Clearly, $\mathbf{P}_{i+1}|_{\alpha_{i}(\mathbf{A})}$ also satisfies Invariant 1 as it has homomorphism-embedding to $\mathbf{A}$ where $\ll_{\mathbf{A}}$ is a partial order and $_{\mathbf{A}}$ its linear extension. It remains to notice that Invariant 1 is preserved by free amalgamations.
Note that $_{\mathbf{P}}$ has a homomorphism-embedding to $_{\mathbf{D}}$ which is a linear order. Consequently, $_{\mathbf{P}}$ is acyclic. The first point of Invariant 1 implies that $\ll_{\mathbf{P}}$ is also acyclic. It remains to construct an $L$-structure $\mathbf{C}$ by putting $C=P$, letting $_{\mathbf{C}}$ extend $_{\mathbf{P}}$ to a linear order, and putting $\ll_{\mathbf{C}}$ to be the transitive closure of $\ll_{\mathbf{P}}$. By the second point of Invariant 1 we know that every embedding $f\colon\mathbf{B}\to\mathbf{P}|_{L}$ is also an embedding $f\colon\mathbf{B}\to\mathbf{C}$. ∎